A solution containg 1 gm per dm3 of urea is isotonic with 5% solution of a nonvolatile solute. What is the molecular mass of this non volatile solute

Grade:12

1 Answers

Vikas TU
6 years ago
Given information : 
Urea arrangement is 1gm for every dm3, so m=1g and V = 1000 ml 
Non unstable arrangement is 5%, so m=5g and V = 100ml 
General Gas condition 
PV = nRT      PV =m/M RT 
PM = m/V RT 
For urea P1 = (m_1 RT)/(M_1 V_1 ) 
For Non unstable Solute P2 = (m_2 RT)/(M_2 V_2 ) 
Since both arrangements are isotonic, so P1 = P2 
Subsequently we can compose 
(m_1 RT)/(M_1 V_1 )=(m_2 RT)/(M_2 V_2 ) 
m_1/(M_1 V_1 )=m_2/(M_2 V_2 ) 
1g/(60 g/mol X 1000ml)=5g/(M_2 X100ml) 
M_2=3000 g/mol 
In this way, atomic weight of non unstable solute is 3000 g/mol.